1. A student takes data on magnetic field as a function of distance for a permanent magnet and fit it to the equation B=A/d^3. They find the A parameter to be 3.87x10-4. What is the value of the magnetic moment (μ)? Answer in A*m2.

2. You have a loop of wire that you want to run a current through to produce an electromagnet. The radius of the loop is 9 mm and the magnetic moment you desire is 1.24 A*m2. What current is required to produce this magnetic moment? Answer in Amps.
Answer #1

1.

The magnetic field at a point on the axis of a bar magnet is B = 10-7 x 2 μ/d3

The given equation for magnetic field is B = A/d3

By comparing the two equations, A = 10-7 x 2 μ

Magnetic moment, μ = A /10-7 x 2 = 3.87 x 10-4/ 10-7 x 2 = 1.94 x 103 Am2

Magnetic moment, μ = 1.94 x 103 Am2

2.

Area of the loop, A = πr2 = 3.14 x ( 9 x 10-3)2 = 2.54 x 10-4 m2

Magnetic moment of a current loop, μ = I A

Current required, I = μ/A = 1.24/2.54 x 10-4 = 4.88 x 103 A

Current required, I = 4.88 x 103 A